• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Sirish

    Log entry time 13:33:44 on July27,2005

    Entry number 149673

    keyword=compton

    Did a vertical beam scan to establish new CIP. (Compton Interation Point). new position IMB102B(V)=-0.75 mm.
    Substantial improvement in signal rate after that. See below.

    However, compton DAQ conicidence trigger hardware has failed. Need an access to troubleshoot. No compton runs till problem is resolved.


    FIGURE 1